# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ A.3 Relation with Recovering All Information
$\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
@@
$\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
$=$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\log{\sum_{j=1}^{N}e^{d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$ (15)
@@
A.4 The Proof of identifiability on hypersphere
A.4.1 The Proof of Theorem 3.1
@@ A.4 The Proof of identifiability on hypersphere
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eqs. (4)-(5) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized symmetric contrastive loss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x}),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t}),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}$ (18)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)},$ (19)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})e^{(\mathbf{h}(\mathbf{z}_{t})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$ (20)
@@
The proof of Theorem 3.1 relies on establishing the equality between the right-hand side of Eq. (18) and that of Eq. (14). Our initial step demonstrates that the global minimum of (14) necessitates $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$ for all real pairs $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$. The second step involves demonstrating that the conditional distribution defined in (4), along with the aforementioned condition in the first step, guarantees two key properties: 1) $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and 2) their independence from the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. Finally, by denoting $\mathbf{h}=\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$ and applying the generative model defined in Eqs. (4)-(5) and the inference model outlined in Eqs (19)-(20), we arrive at the conclusion of the theorem.
@@
On hypersphere space, the $d$ distance metric defined as cosine similarity in Eq. (14), thus we can rewrite the right-hand side of Eq. (14) as follows:
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{x},\mathbf{t})\sim p(\mathbf{x},\mathbf{t})}{\big{[}\big{(}\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{\big{(}\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{\big{(}\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}.$ (21)
Utilizing the generative model described in Eqs. (4)-(5), we can reframe Eq. (21) in terms of the latent variables $(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})$. Furthermore, by introducing the notations $\mathbf{h}_{x}=\mathbf{f}_{x}\circ\mathbf{g}_{x}$ and $\mathbf{h}_{t}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, we can express Eq. (21) as follows:
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau\big{]}}$
$+$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{m}_{x},\mathbf{z}_{x})\sim p(\mathbf{m}_{x},\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{(\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{t},\mathbf{z}_{t})}\big{[}e^{\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$
$+\operatorname*{\mathbb{E}}\limits_{(\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{t},\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{(\mathbf{m}_{x},\mathbf{z}_{x})\sim p(\mathbf{m}_{t},\mathbf{z}_{t})}\big{[}e^{\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (22)
Further, since we assume the inference model as define in Eqs (19)-(20), and $\tau$ can be optimized in multimodel contrastive representation learning, we can further rewrite Eq. (22) as follows (see Appendix A.3 for details):
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau\big{]}}$
$-H(p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})),p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})))+\log Z_{\text{KDE}}$
$-H(p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})),p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})))+\log Z_{\text{KDE}}.$ (23)
This can be further formulated using the Kullback–Leibler divergence as follows:
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau\big{]}}$ (24)
$-H(p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})))-D_{\text{KL}}(p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x}))||p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})))+\log Z_{\text{KDE}}$ (25)
$-H(p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})))-D_{\text{KL}}(p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t}))||p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})))+\log Z_{\text{KDE}}.$ (26)
On hypersphere space with radius $r$, due to $\lVert\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})-\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\rVert=2r-2\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})^{T}\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$, the first term (24) is minimized if and only if
[…]
In addition, the last two terms, (25) and (26) is minimized if both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map $(\mathbf{m}_{x},\mathbf{z}_{x})$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$, respectively, to uniform variables on hypersphere (condition 2, note that here uniform distributions also supports the symmetry of Eq. (19) and Eq. (20)). Since condition 1 does not affect the validity of condition 2, it follows that condition 1 is a necessary condition for minimizing the right-hand side of Eq. (18).
@@
By differentiating the equation with respect to $\mathbf{m}_{x}$ in the condition 1 in the step I, we have:
, due to the independence between $\mathbf{m}_{x}$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$. Similarly, by differentiating the equation in the condition 1 with respect to $\mathbf{m}_{t}$, we have:
Based on Eqs. (28) and (29), we conclude that both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, i.e., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$ and $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{t})$. Further, since $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{t})$ hold, for all real pairs $(\mathbf{z}_{x},\mathbf{z}_{t})$ sampled from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ defined in Eq. (4), this expression also holds true for $\mathbf{z}_{t}=\mathbf{z}_{x}$, which implies $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{x})$. As a result, we can obtain: $\mathbf{h}_{x}=\mathbf{h}_{t}$.
Step III
@@
According to the results above: $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$, $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{x})$, and $\mathbf{h}_{x}=\mathbf{h}_{t}$ from Step II, by defining $\mathbf{h}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbf{h}_{x}=\mathbf{h}_{t}$ , we can rewrite Eq. (21) as:
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (30)

### v2

@@ Appendix C Relation with Recovering All Information
$\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}+\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
@@
$\operatorname*{\mathbb{E}}\limits_{\mathbf{x}\sim p(\mathbf{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{t}\sim p(\mathbf{t})}\big{[}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}),\mathbf{f}_{t}(\mathbf{t})\big{)}/\tau}\big{]}\Big{]}$
$=$ $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{N}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}}$ (11)
@@
Appendix D The Proof of identifiability on hypersphere
D.1 The Proof of Theorem 4.1
@@ Appendix D The Proof of identifiability on hypersphere
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eq. (4) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized multimodal contrastiveloss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x}),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t}),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}$ (14)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})^{-1}e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)},$ (15)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})^{-1}e^{(\mathbf{h}(\mathbf{z}_{t})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$ (16)
@@
Lemma 5.1.
The normalized multimodal contrastive loss in Eq. (2) has an optimal global solution of 0, which can be attained under the following conditions:
  • $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ almost surely, for pair $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$,  (C1),
  • $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map $(\mathbf{m}_{x},\mathbf{z}_{x})$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$, respectively, to uniform variables on hypersphere, (C2),
Proof.
First, it is well known that traditional contrastive loss in single modality has an optimal global solution of $\log N$ [43, 56], as a result, the multimodal contrastive loss Eq. 1 has an optimal global solution of $2\log N$. For completeness, let us focus on the first term in Eq. 1:
$-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}}{\sum_{j=1}^{N}e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}},$ (17)
Under optimal contrastive learning conditions, the distance for positive pairs satisfies: $e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{i})\big{)}/\tau}=1$, for negative pairs $(\mathbf{x}_{i},\mathbf{x}_{j})$ where $i\neq j$: $e^{-d\big{(}\mathbf{f}_{x}(\mathbf{x}_{i}),\mathbf{f}_{t}(\mathbf{t}_{j})\big{)}/\tau}=\epsilon$, where $\epsilon$ is a small value. As a result, for each $i$, the denominator can be expressed as: $1+(N-1)\epsilon$. Therefore, the first term in Eq. 1 reduces to : $-\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{1+(N-1)\epsilon}$. Clearly, when $N$ is large, the first term in Eq. 1 equals to $\log N$. Given that the second term is symmetric, we conclude that Eq. 1 has an optimal global solution of $2\log N$. Therefore, Eq. 10 achieves a global optimal solution of 0. To reach the global minimum of 0, we observe that the first term in Eq. 10 is minimized if and only if $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ almost surely, for real pair $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$, (marked as (C1)). Thus, we obtain a minimum solution of 0 for the first term. Next, considering the remaining two terms in Eq. 10, as detailed in Appendix C, we see an equivalent expression: $-H(p(\mathbf{f}_{x}(\mathbf{x}),p(\mathbf{f}_{t}(\mathbf{t})))-H(p(\mathbf{f}_{x}(\mathbf{x}),p(\mathbf{f}_{t}(\mathbf{t})))+2\log Z_{\text{KDE}}$. When both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map $(\mathbf{m}_{x},\mathbf{z}_{x})$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$, respectively, to uniform variables on hypersphere (marked as (C2)), it reduces to $-2H(p(\mathbf{f}_{x}(\mathbf{x}))+2\log Z_{\text{KDE}}$. Note that the entropy of a uniform distribution on the hypersphere $\mathbb{S}^{M-1}$ is $\log(\frac{2\pi^{M/2}}{\Gamma(M/2)})$, where $\Gamma$ is the gamma function. Together with the fact that the normalization constant of uniform distribution on hypersphere is $\log(\frac{2\pi^{M/2}}{\Gamma(M/2)})$ (i.e., $\log Z_{\text{KDE}}$), we arrive at the optimal solution of 0 for the last two terms. ∎
Proof sketch
The proof of Theorem 4.1 hinges on demonstrating the equality between the right-hand side of Eq. (14) and Eq. (10). Let us define $\mathbf{h}_{x}=\mathbf{f}_{x}\circ\mathbf{g}_{x}$ and $\mathbf{h}_{t}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$. In Step I, using Lemma 5.1, we show that (1) $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and (2) they are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$. In Step II, by defining $\mathbf{h}=\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$ and applying both the generative model from Eq. (4) and the inference model from Eqs. (15)-(16), we establish the theorem.
@@
Consider C1 in Lemma 5.1, e.g., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ almost surely, for pair $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$, by differentiating it with respect to $\mathbf{m}_{x}$, we have:
, due to the independence between $\mathbf{m}_{x}$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$. Similarly, by differentiating it with respect to $\mathbf{m}_{t}$, we have:
Based on Eqs. (18) and (19), we conclude that both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, i.e., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$ and $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{t})$. As a result, we have $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{t})$, for all real pairs $(\mathbf{z}_{x},\mathbf{z}_{t})$ sampled from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ defined in Eq. (4). Note that this expression also holds true for $\mathbf{z}_{t}=\mathbf{z}_{x}$ (e.g., when $\mathbf{z}_{t}$ is sampled with the same value as $\mathbf{z}_{x}$), which implies $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{x})$. As a result, we can obtain: $\mathbf{h}_{x}=\mathbf{h}_{t}$.
@@
According to the results above: $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$, $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{x})$, and $\mathbf{h}_{x}=\mathbf{h}_{t}$ from Step I, by defining $\mathbf{h}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbf{h}_{x}=\mathbf{h}_{t}$ , we can rewrite Eq. (10) as:
$2\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}$ $+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (20)